6/17/99 Day Shift

Yee Bob Hsiung (hsiung@fnal.gov)
Thu, 17 Jun 1999 15:47:39 -0500 (CDT)

Shift Summary for DAY shift, 17-June-99

On Shift: Yee Bob Hsiung
Emmanuel Monnier

Hours of beam delivered at all intensities: 6h
Hours of beam delivered at nominal intensity: 5h
Hours of E832 data (special runs not included): 5h
Hours of special runs (laser scans, muon runs...): 0h

Accounting of lost time:

Unstable beam from time to time plus problems of QXR
and TeV F2 VCB opened.

========================================================================
Runs ended this shift:

13634 Beam_832 13.6M events, 20.4k Vac00 <SEM>=3.7E12
13635 Beam_832 0.29M events. on new tapes
13638 Beam_832 in progress 4M so far.

Tapes brought to FCC this shift:
QKB005
QKC013,QKC014,QKC015
QKK013,QKK014,QKK015
QKS013,QKS014,QKS015
(for runs 13624-13634)
========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

Much smooth data taking with from time to time sporadic beam.
Fixed hang up problem when stopping the run.
2 Ped exponent >0 alarms from Val, but disappeared right next spill.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13634 9:47 lost beam, QXR problem.
10:20 Run badly stoped because mon_alr hang up
forget to save the histogram for this long run.
Made plots for online reference plot binder
from this run.
10:34 Ed updated CSI constants valid for run 13635 & up
10:45 beam back
13635 10:50 Start run with new tapes
11:14 Beam stopped because of TeV F2 VCB openned
11:30 Run stopped to solved mon_alr problem -> solved.
ECC remake the L3driver and monitoring code.
Several drt_fresh_starts.
Noticed tape labeling takes 5-6min at the start
of the run.
13638 12:00 Start new run just in time with return of beam.

14:35 CO2 pressure for DC very low <200psi (bottom of
the scale) switch to new bottles. now 900psi.

========================================================================
Follow up Notes from Previous Shift

Red bottons for 3 tape drives for online split due to empty drive
was allocated. Sydney fixed the problem and is investigating
why it happens (i.e. Was shiftee ignoring the tape mount requests
and the process timed out ~2hrs?)

=======================================================================
Items to Follow up on Next Shift

NM2EU1 unstable according to Rick Ford. He need 2hr to fix it, if there
is downtime.
Tomorrow morning after 9am there will be radiation survey in KTeV enclosure
for 1-2 hours. (We maybe able to run some muons with sweeper magnet set
high)